What is scattering amplitude in born approximation.
1
Expert's answer
2020-11-11T08:02:57-0500

Solution

For born approximation scattering amplitude is given by

"f=-\frac{2\mu }{\hbar^2 k}\int rV(r) \sin kr dr"

Valid for central potential.

Where

"\mu" is reduced mass

"\hbar" is reduced planck constant

V(r) is central potential.